OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12

OpenStudy (aripotta):

do you know what slope-intercept form is?

OpenStudy (anonymous):

\[9y = 11x -12\] \[y = (11x-12)/9\] \[y= 11x/9-12/9\] We're solving for y

OpenStudy (anonymous):

unfortuenly no

OpenStudy (aripotta):

slope-intercept form is y = mx + b. where m is the slope and b is the y-intercept

OpenStudy (anonymous):

slop is 11/9

OpenStudy (anonymous):

nvm i confusing

OpenStudy (anonymous):

Its in Standard Form. Ax + By = C You need it in slope-intercept form. y = mx + b. Basically you need to isolate y on one side. SO: -11x + 9y = -12 Add 11x to both sides. 9y = 11x - 12 Divide both sides by 9. y = 11/9x - 12

OpenStudy (anonymous):

Whoops I forgot to divide 12 by 9. SO it would be: y = 11/9x - 12/9 Or as a decimal... y = 1.22x - 1.33

OpenStudy (anonymous):

i dont undestand any of this

OpenStudy (anonymous):

its a or c

OpenStudy (aripotta):

ok, take a deep breath, close your eyes, and think of nothing for 30 seconds.

OpenStudy (aripotta):

so you can calm down and return with a clear and open mind

OpenStudy (anonymous):

thats funny

OpenStudy (aripotta):

it's not supposed to be funny.

OpenStudy (aripotta):

you can't keep relying on people to be there to "check you work" so you need to learn how to do things

OpenStudy (aripotta):

your*

OpenStudy (anonymous):

so it is not a or c or is it

OpenStudy (aripotta):

i don't know, you didn't post the choices

OpenStudy (anonymous):

lol its c

OpenStudy (aripotta):

can you post it so i can make sure for you?

OpenStudy (aripotta):

or did you just turn it in? :/

OpenStudy (anonymous):

im on next question